OpenStudy (anonymous):

Is the translation for g(x)=-(2^1/2(x+3))-5 is: Reflection in the x-axis, Horizontal stretch by a factor of 2, horizontal translation 3 units left, vertical translation 5 units down. Is this correct?

OpenStudy (xishem):

It depends on what the original function was. Also, unclear about what is written. Is it this? \[g(x)=-(\frac{2^1}{2(x+3)}-5)\]

OpenStudy (xishem):

\[g(x)=-(\frac{2^1}{2(x+3)})-5\]Sorry, I meant that.

OpenStudy (anonymous):

\[g(x)=-(2^{\frac{ 1 }{ 2 }(x+3)} )-5\]

OpenStudy (anonymous):

this is the equation